Report of Activities on the RVIB N.B. Palmer Cruise 02-02
4 May 2002
On 4 May, the N.B.
Palmer was working along the middle of the continental shelf on SO GLOBEC
survey line 9. The seas remained
moderate. The clouds were thicker than
yesterday, but higher and the visibility was good. The mountains of
Just after 1600, the fire alarm went off. This time it was a drill. Within a few minutes all in the scientific party had appeared at the muster station ready, if necessary to abandon ship. There were quite a few sleepy faces of those on the watch who had been awakened by the alarm. The drill ended with everyone signing the bridge book before leaving the 03 level lounge.
Work was completed at broad-scale survey stations 72, 73, and 74 including 3 CTDs, a MOCNESS tow at station 73, and a 1-m ring net tow and an APOP cast at station 74. BIOMAPER-II was in for transits only a portion of the stations because of a ground fault problem with the Environmental Sensing System. Seabird and marine mammal surveys took place during the daylight period and two sonobuoys were deployed during the transit to station 74.
CTD Group report (John Klinck, Tim Boyer, Chris Mackay, Julian Ashford, Andres Sepulveda, Kristin Cobb)
The CTD group did three CTD casts across the shelf on the 180 line (survey line 9). The two deeper stations were typical of the shelf with deep uniform mixed layers and bottom temperatures of 1.3ºC. The shallower station near the coast had a thinner mixed layer and a weaker pycnocline.
Station 72 (cast 75, 438 m). There was a uniform mixed layer to 65 m (-1.7ºC, 33.4psu , 0.1 ug/l chlorophyll) and a relic Winter Water (WW) layer at 90 m. The pycnocline extended to 250 m with thin (3 m) temperature reversals. The deep temperature was uniform (1.3ºC).
Station 73 (cast 76, 523 m). A uniform mixed layer occurred to 55 m (-1.7ºC, 33.4 psu, 0.1 ug/l chlorophyll) and relic WW layer was at 80 m (30 m thick). The pycnocline extended to 250 m with small-scale variability. The deep temperature was uniform (1.3ºC).
Station 74 (cast 77, 224 m). A uniform mixed layer occurred to 35 m (-1.8ºC, 33.4 psu, 0.07 ug/l chlorophyll). Salinity increased slowly below 50 m to the bottom. There was a 20 m thick higher temperature layer (0.7ºC) at the bottom.
Nutrients Status Report (Rob Masserini)
Nutrient analysis continued in the hydrolab on the NBP. Since the last report (27 April), an additional 511 seawater samples from 23 CTD casts have been run and the data processed. This brings the total to 1400 samples to date, culminating in 7000 separate analysis for nitrate, nitrite, ammonia, phosphate, and silicic acid, not to mention an almost absurd number of standards and quality control samples - somewhere on the order of 2250 samples or 11250 analyses.
Preliminary cross-section plots generated by Yulia Serevennikova for transects
7 and 8 indicate that the general hydrographic structure below the mixed layer for
the shelf is consistent with transects 1 through 6 (please see earlier reports
for concentration particulars). The
stations inside of
Marine Mammal report (Debra Glasgow)
Saturday, 4 May dawned overcast and very cold. At 1000, the air temperature was -5.4ºC and there were problems keeping the inside of the windows in the ice tower free of ice. It seemed more time was spent scraping ice than doing survey. With the sea state at Beaufort 5-6, most of the day conditions were borderline for survey. However, survey began anyway while a MOCNESS was deployed at station 73 at 1013 and continued along the transect line towards Station 74. Several icebergs, growlers, and ice bits were past during the day and a total of 3 fur seals were recorded.
At 1423, an anomalous movement was seen in the water 145º, 2.5 nautical miles ahead, and using the 7x20 binoculars, one humpback was seen to surface twice. The first time only the body was seen. On the second, it was the body and the blow. It was swimming towards 040º and was recorded at 68 32.10ºS; 73 29.64ºW as WOS#24. The first fur seal of the day was recorded 5 minutes later. At 1557, a black body was sighted 4.5 nautical miles away, bearing 100º ahead at 68 36.42ºS; 73 12.74ºW and confirmed that this was one of a group of at least four humpbacks on the edge of grease ice/shuga. This whale part breached from its head to the pectoral fins out of the water, and then continued to wave and slap its pectoral fins creating a lot of splashes. The other whales nearby seemed to lunge and be active also, but not as much as the first whale described, two of them were mostly just resting on the surface. This was sighting number WOS#25.
There were two other distinct groups of humpbacks nearby - WOS#26 sighting of 2+ humpbacks at 68 37.03ºS; 73 10.78ºW, 6 nautical miles away, 070º, swimming to 150º, and WOS#27 sighting at 68 37.29ºS; 73 09.91ºW of another pair of humpbacks also causing splashes at approximately 6 nautical miles away at 110º. Unfortunately, the fire alarm went off at this point for what turned out to be a drill, so I could not track these groups or get a better estimate of numbers or activity. All these whales were within the same general area close to the edge, but within an extensive area of grease ice/shuga.
Seconds before the fire alarm bell went off at 1617, a humpback surfaced directly beside the port bow of the ship (within 30 meters) at 120º, 68 37.56ºS; 73 09.06ºW. It was swimming very fast alongside the ship heading to 130º. This was recorded as sighting WOS#28. When I returned to the ice tower at 1632, the whales had disappeared and the light was fading fast so I was unable to resight these whales. Survey ended in very poor light at 1643. A total of 10 humpback whales were recorded today.
Sea Birds (Erik Chapman and Matthew Becker)
There were approximately 40 Snow Petrels circling the ship
when the Palmer left station 73 this
morning (4 May) and headed toward
A surface net tow was conducted at station 74, producing no diatoms and a few larval krill and larval fish. Larval krill and fish are representative of under-ice biology and are likely prey that Snow Petrels typically focus on as they forage along the interface between open water and ice edge. Station 74 was characterized by 8/10ths ice coverage of small, 2 m diameter floes, covered in about 50 cm of snow and some brash.
During the fall cruise last year, a large number of Snow Petrels were seen in the same survey area, both in open water and ice habitats. This area near Alexander Island appears to be a focal point for early winter ice production and seems to generally be characterized by high ice coverage during the early fall. Due to the strong association of Snow Petrels with ice coverage, such conditions would presumably prove very attractive to these birds. Proximity to ice coverage throughout much of the year, including the fall and presumably the early spring, may also explain why Alexander Island is the site of a relatively large number of Snow Petrel breeding sites on nunataks (exposed rock outcropping through glacial ice) and cliff faces.
Snow Petrels may remain in these areas after breeding until
the winter pack-ice develops, after which they radiate out into the ocean to
work open water in the pack ice and ice edges.
Indeed, this is a pattern that was observed as Snow Petrel distribution
shifted dramatically from the fall to the winter cruise last year. During the winter cruise, the study grid was
covered in sea-ice and Snow Petrels were no longer primarily concentrated near
Night observations were performed in the late evening under
calm, clear skies, and 10/10 ice coverage.
Bird activity was reduced to a single snow petrel observation, with no
other species sighted. No penguins or
penguin tracks were visible on the ice.
Braving the cold, we were rewarded with a spectacular stars and a moon
rise over the mountains of
A summary of the species and number of individuals of birds and seals within the 300 m transect during 3 hours, 23 minutes of daytime surveys between consecutive stations 73 and 74 is the following:
Species (common name)
Species (scientific name)
Southern Giant Petrel
Material Properties of Zooplankton Report (Dezang Chu, Peter Wiebe)
On May 4, measurements of material properties of E. superba
were made at station 74 (68 40.931ºS, 72 54.889ºW), northwest of the
Shortly before the APOP cast, a loose connection was found between the wire and ceramic disk in one of the transducers, occasionally reducing the received signal level and altering the phase, significantly. Since we didn’t have enough time to fix the problem, we decided to deploy the APOP anyway. We tied the transducer cables to the APOP framed firmly, hoping they would not swing around during the measurements. However, at the moment the APOP entered the water, a severe rattling motion of the APOP was observed and signal level of one of the two channels was reduced. Since we observed consistent signals in amplitude and phase after the first disturbance, we decided to continue the cast without pulling the APOP. The whole cast was, in general, went well, except that in the middle of the up cast, there was another observed phase jump. After we compensated the phase changes due to bad cable connections, we actually obtained very good results. There was basically no difference between the down and up casts. The difference was only 0.0005, about of an order of magnitude lower than some of the previous casts. The mean and standard deviation of the sound speed contrasts were 1.021 and 0.001, respectively. The variation was very small, indicating consistent measurements. The sound speed contrast was nearly constant over 200 meters, but a tendency of reduction in sound speed contrast with depth was also noticeable, 0.01 to 0.02 over 200 meters. The mean value almost exactly the same as that we obtained for the same species of similar size animals used in a previous APOP cast (see report on April 27).
The density measurement was not completed because of a broken syringe. The mixed solution was split out before it was well mixed. We had to measure the displacement volume to estimate the volume of the animals with a graduated cylinder.
Zooplankton (MOCNESS/BIOMAPER-II) report (Carin Ashjian, Peter Wiebe)
MOCNESS Tow #18 was conducted at mid-shelf station 73 along Transect #9, concluding just after and sampling down to 525 m. Low abundances of plankton were observed throughout the water column. Copepods were the dominant taxa, being found at all depths. The bioluminescent copepod Metridia was especially obvious in the 25-150 m intervals. A large, purple jellyfish was collected in the deepest net (350-525 m). Chaetognaths were found in the deeper nets from 150 - 525 m. Very small larval krill were collected at depth as well, from 150 - 525 m, but were not abundant. Larval krill also were observed in scant abundances in the upper 50 m. Amphipods were seen at depth (250-525 m) and at the surface (0-25 m).
The along track survey of zooplankton and nekton using BIOMAPER-II was sidelined for a portion of the transits between stations 71, 72 and 73. The cause was due to the failure of the Environmental Sensing System (ESS) to turn on once the towed body was in the water. During the previous transit, an electrical fault indicator alarm had been intermittently going on and then off, but checks of the underwater connectors, the usual cause of such a fault, showed no problems. It was ultimately determined to be due to a water leak in the ESSs pressure sensor tubing in underwater electronics unit. Fortunately, although a half-cup of water was in the case, there was no damage. Cleaning of the circuitry with alcohol and contact cleaner, and refitting of the pressure unit tubing by Andy Girard, put the unit back in service.
Steaming towards station 73 in the mid-continental shelf region with bottom depths over 600 m, there was low volume backscattering in the upper portion of the water column on all echograms and a bit more scattering deep that was probably associated with the bottom, although the bottom was too deep to be seen on the echogram. On way to station 74, there was light scattering between 100 meters and 180 meters, while the sea surface had no ice. After the ship entered a region of grease and shuga ice, a thick layer of scattering between 75 and 120 m developed and dense patches were seen at 40 m. In the evening, upon leaving station 74 on the run towards station 75, there was very strong krill-like backscattering just below the surface, starting about 40 m. The VPR images showed the layer to have both copepods and krill in it. The strong back scattering diminished after several miles and the backscattering, although moderately strong, was not krill-like. Along this trackline, the bottom shoaled and became highly variable resembling a roller coaster ride. This rapid variation in bottom depths made towyoing BIOMAPER-II to any depth nerve racking. Before , a very strong backscattering layer developed right above the bottom, which by this time was quite shallow (around 146 m). The layer, which was about 20 m thick, only persisted for a short time and them the scattering decreased substantially.